OpenStudy (anonymous):

whats the y intercept and x intercept of y=-3x+2 show all your work

OpenStudy (anonymous):

y int. is 2 (just look at the equation) and x int. is: y=0 0= -3x+2 -2=-3x 2/3 = x x int. is 2/3 (:

OpenStudy (slaaibak):

y intercept -> make x zero: y=-3(0)+2 y=2 x intercept -> make y zero: 0=-3x+2 3x=2 x=2/3